# Electromagnetic Radiation from High-Energy Nuclear Collisions

### v1

@@ 1 Introduction
The detection of the electromagnetic radiation emanating from the collision of strongly interacting particles dates back to the very early scattering experiments and even predates the establishment of the theory of the nuclear strong interaction, QCD. Owing to the relative weakness of the electromagnetic coupling, $\alpha_{\rm EM}/\alpha_{\rm s}\ll 1$, photons (real and virtual) offered the promise of a clean probe of the reaction dynamics as the mean free paths associated with the electromagnetic interaction exceed typical hadronic length scales. In the modern era ushered by the availability of high energy hadronic beams, the immense potential of electromagnetic observables became clear, and their status as natural complements to measurements of hadrons was confidently established. To make the discussion specific, the photon production cross section in proton-proton collisions, for example, is calculable in perturbative QCD (pQCD), with the help of factorization Field:1989uq . Consequently, early measurements of electromagnetic radiation at CERN (ISR) and at Fermilab where crucial in the confirmation of the parton model, of QCD itself, and of its perturbative treatment Drell:1970wh ; Owens:1986mp .
To repeat, the basic feature which makes real and virtual photons so appealing in studies of QCD is that they suffer no final state interactions and decouple from the strongly interacting system. They are therefore good probes of the conditions local to their production site, assuming their formation is under theoretical control. This is also the reason why electromagnetic measurements have accompanied the development of the field now known as relativistic heavy-ion collisions, which concerns itself with the formation, study, and characterization of the quark-gluon plasma, an exotic state of the fundamental constituents of QCD and which is also the central theme of the present volume.
@@
Figure 1: The many stages of a heavy-ion collision: (a) Initial state; (b) pre-hydro phase; (c) fluid dynamical evolution; (d) Hadronization; (e) Final-state hadrons eventually free-streaming to the detectors. These images are not from a single end-to-end calculation, but are simply meant to illustrate the different stages. The pictures have been assembled and collected from different sources which include hydrodynamical simulations Schenke:2010nt and UrQMD calculations Bleicher:1999xi .
@@ 2 Electromagnetic radiation from systems in equilibrium
Some of the early exploratory calculations have been devoted to estimates of the temperature of the quark-gluon plasma, and have proposed measuring the spectrum of lepton pairs with invariant masses $1<M\lesssim 3.5$ GeV, and read an effective temperature from its slope in analogy with black body radiation Shuryak:1978ij ; Kajantie:1981wg ; Domokos:1980ba ; Kajantie:1986dh . This basic philosophy – using electromagnetic radiation to characterize the QGP – is followed to this day, but with some of the additional sophistication described in this paper.
@@
In a strongly interacting medium at temperature $T=1/\beta$, a starting point to the computation of thermal electromagnetic emission rates is the transition rate between two states, an initial state $i$ and a final state $f$, which can be written as $R_{fi}=\frac{\left|S_{fi}\right|^{2}}{\tau V}$ with $\tau V$ being the proper four-volume. To leading order in the electromagnetic interaction, the interaction matrix element written in terms of the electromagnetic current $J^{\mu}$ and the photon field $A^{\mu}$ is $S_{fi}=\langle f|\int\!\!d^{4}xJ^{\mu}(x)A_{\mu}(x)|i\rangle$. A by now standard set of manipulations Feinberg:1976ua ; McLerran:1984ay ; Weldon:1990iw ; Gale:1990pn ; Kapusta:2007xjq lead to the expressions for the differential emission rate $R$ (electromagnetic quanta per unit four-volume) of photons and dileptons:
$\omega\frac{dR_{\gamma}}{d^{4}k}$ $=$ $-\frac{g^{\mu\nu}}{\left(2\pi\right)^{3}}{\rm Im}\,\Pi_{\mu\nu}(\omega,{\bf k})\frac{1}{e^{\beta\omega}-1}\ \ {\rm(photons)}$
$E_{+}E_{-}\frac{dR_{\ell\bar{\ell}}}{d^{3}p_{+}d^{3}p_{-}}$ $=$ $\frac{2e^{2}}{\left(2\pi\right)^{6}}\frac{1}{k^{4}}L^{\mu\nu}\,{\rm Im}\Pi_{\mu\nu}(\omega,{\bf k})\frac{1}{e^{\beta\omega}-1}\ \ {\rm(dileptons)}$ (1)
@@
The lepton tensor is (for leptons with rest mass $m_{\ell}$) $L^{\mu\nu}=p_{+}^{\mu}p_{-}^{\nu}+p_{+}^{\nu}p_{-}^{\mu}-g^{\mu\nu}\left(p_{+}\cdot p_{-}+m_{\ell}^{2}\right)$. The photon (dilepton) four-momentum $k^{\mu}=(\omega,{\bf k})$ is light-like for real photons and timelike for dileptons. In these equations, $\Pi^{\mu\nu}$ is the finite-temperature retarded in-medium photon self-energy. Importantly, these equations are valid up to first order in the electromagnetic coupling, but are exact in the strong coupling Gale:1990pn . This formalism is then directly amenable to non-perturbative field theoretical calculations, where the photon self-energy can be written as a two-point correlator of the electromagnetic current Gale:1990pn obtainable by analytically continuing the Euclidean correlator evaluated non-perturbatively on the lattice. For a non-Abelian gauge theory like QCD this is a technically challenging procedure Ghiglieri:2016tvj ; Ce:2022fot , but recent progress has been reported Ali:2024xae .
@@
An alternative approach to the evaluation of electromagnetic radiation production rates is provided by the fact that taking the imaginary part of the photon self-energy is tantamount to putting internal virtual particles on-shell, enabling in principle a formalism based on kinetic theory Weldon:1983jn ; McLerran:1984ay ; Majumder:2001iy . In this language, the photon production rate from the reaction $1+2\to 3+\gamma$ is
$R_{\gamma}$ $=$ ${\cal N}\int\frac{d^{3}p_{1}}{2E_{1}\left(2\pi\right)^{3}}\frac{d^{3}p_{2}}{2E_{2}\left(2\pi\right)^{3}}f_{1}(E_{1})f_{2}(E_{2})\left(2\pi\right)^{4}\delta\left(p_{1}^{\mu}+p_{2}^{\mu}-p_{3}^{\mu}-k^{\mu}\right)$ (2)
×$\times\left|{\cal M}\right|^{2}\frac{d^{3}p_{3}}{2E_{3}\left(2\pi\right)^{3}}\frac{d^{3}k}{2\omega\left(2\pi\right)^{3}}\left[1\pm f_{3}(E_{3})\right]$
where ${\cal N}$ is an overall degeneracy factor, the $f$’s are Fermi-Dirac or Bose-Einstein distribution functions as appropriate, and the matrix element for this interaction is ${\cal M}$. Those two approaches have their own advantages and disadvantages, and are often used as being complementary.
@@
The historical development of thermal emission rates from strongly interacting matter as that created in heavy-ion collisions is an interesting story in itself. Calculations of photons and dileptons from the hot QGP involve the interactions between gluons and quarks in a finite-temperature medium. Early kinetic theory calculations provided estimates of photon and dilepton emission from interacting partons Shuryak:1978ij ; Kajantie:1981wg ; Sinha:1983jm ; Hwa:1985xg ; Staadt:1985uc . Reformulating the production rate of real photons in terms of the photon self-energy with Eq. (1) in mind, the processes at leading order in $\alpha_{s}$: $qg\to q\gamma$ and $q\bar{q}\to g\gamma$ appear at two loops. If the virtual quarks exchanged in those previous reactions are massless, the emission rate suffers from a logarithmic collinear singularity, cured however by the generation of an in-medium effective fermion mass calculated using the Hard Thermal Loops (HTL) resummation Kapusta:1991qp ; Baier:1991em . The formalism of HTL was established earlier and used to famously address the manifest and vexing gauge dependence that plagued calculations of the gluon damping rate in finite temperature QCD [^1] See Refs. [3] in Ref. Braaten:1989mz ..
@@
Figure 2: Two photon-producing processes that come from taking the imaginary part of the two-loop photon self-energy. They are (left) bremsstrahlung from a quark (or anti-quark), and (right) $q{\bar{q}}$ annihilation with scattering.
@@
Further explorations in the topology of the photon self-energy lead to evaluations of the parton bremsstrahlung contribution to the production of real photons. Using an effective theory based on the resummation of HTLs, this channel enters at the two-loop level, and taking the imaginary parts produces the processes shown in Fig. 2. Even though the naive parametric power counting of the photon self-energy would have those contributions $\sim g^{4}$, where $g$ is the strong coupling constant, it was found that collinear divergences formally promote the photon rate to $\sim g^{2}$ Aurenche:1998nw . The solution to the puzzle of how to uncover and sum up all the contributions at a given order of the perturbation expansion was provided with the consistent inclusion of photons generated by the Landau-Pomeranchuk-Migdal effect (LPM), which limits the coherence length of the emitted radiation. The first photon production rates complete at leading order (LO) in the strong coupling were presented in Ref. Arnold:2001ms , and results at NLO were reported later in Ref. Ghiglieri:2013gia  [^2] For values of $\alpha_{s}$ relevant for the phenomenology discussed here, NLO + LO results show an enhancement of $\sim 20\%$ over LO, owing to some partial cancellations in the NLO contributions..
@@
At lower temperatures, the photon rates calculated with the hadronic degrees of freedom relevant to stages (d) and (e) of Fig. 1, were calculated using interactions modeled with effective chiral hadronic Lagrangians in Refs. Meissner:1987ge ; Kapusta:1991qp ; Song:1993ae ; Turbide:2003si ; Heffernan:2014mla ; Holt:2015cda . Calculations of dilepton production involving composite hadrons also rely on Eq. (1), but the virtual nature of the photon field reveals the vector meson spectral distribution, owing to the coupling of the photon field with vector meson fields: Vector Meson Dominance Gale:1990pn ; Rapp:1999ej .
It is important to appreciate the perseverance and devotion of the experimental community involved in the measurement of photons and dileptons. Measuring lepton pair production is especially challenging: the rates are smaller than their photon counterpart by a factor of $\alpha_{\rm EM}$, and the analysis requires extraction of the desired signal from a huge combinatorial background.
@@ 2.1 Electromagnetic emission out of equilibrium
To fix ideas, we first consider the emission of real photons; of course, virtual photons are also emitted out of equilibrium. Before discussing the different models addressing photon emission out of equilibrium and just prior to the hydrodynamics phase, it is instructive to discuss the most extreme case: electromagnetic radiation from the very first nucleon-nucleon collisions. Those photons will appear as an irreducible background to any of the other photon sources under consideration, and will typically occupy the highest $p_{T}$ part of the spectrum. As alluded to previously, those are calculable with pQCD, using a procedure that is however not free from uncertainties which grow as $p_{T}$ decreases: they are associated with scale uncertainties in the nuclear parton density functions (nPDF) Aurenche:2006vj ; Arleo:2011gc ; Klasen:2013mga ; Paakkinen:2017jpo and with the relatively poorly known photon fragmentation function (especially from gluons) Klasen:2014xfa . This is why the pQCD contributions in the low momentum part $p_{T}~{}\approx~{}1$ GeV/c of the thermal photon window are mostly extrapolations Paquet:2015lta which emphasize the importance of direct measurements of low transverse momentum photons in pp and pA collisions.
@@
Just prior to, but also in the fluid phase, non-equilibrium aspects will manifest themselves in different ways. In the recent couple of decades, the relativistic heavy-ion program pursuing the characterization of the QGP has revealed the remarkable success of viscous fluid dynamical modeling. By comparing with hadronic data, it became apparent that the results of the early calculations based on inviscid hydrodynamics could be brought closer to measurements with the inclusion of transport coefficients such as shear and bulk viscosity Gale:2013da ; the electromagnetic emissivity evaluations also needed to be adapted consequently.
@@
Deviations from kinetic equilibrium can be obtained by considering the linearized Boltzmann equation and evaluating the deviations ($\delta f$) of the distribution functions ($f=f_{0}+\delta f$) from their equilibrium form ($f_{0}$), using techniques Denicol pioneered by Chapman and Enskog Chapman , and by Grad Grad . Using that approach, the photoproduction channels in the QGP and in the confined hadronic sector involving $2\to 2$ kinematics can be corrected for the presence of shear and bulk viscosity Dion:2011pp demanded for consistency with the fluid dynamical evolution HeinzSchenke . The electromagnetic emission rates derived thusly can then be implemented in parallel with the fluid dynamical evolution Paquet:2015lta . Part of this general framework for correcting the photon emission rates is discussed in Ref. Dion:2011pp , and a discussion of the $2\to 2$ photon-producing contributions in the QGP (Compton and quark-anti-quark annihilation) appears in Ref. Shen:2014nfa . To this day not all the photoproduction channels currently include such out-of-equilibrium corrections. In the QGP, this is the case for the photon contribution related to the LPM effect.
@@
It is known that formally, a field-theoretic formulation appropriate for non-equilibrium contributions is based on the real-time formalism as expressed in the “1/2” [^3] The numbers label two branches in the complex time plane Schwinger1961 ; Keldysh:1964ud . Schwinger1961 ; Keldysh:1964ud and “r/a” bases Keldysh:1964ud . In that language, the thermal photon spectrum is Serreau:2003wr
@@
where the photon polarization tensor is no longer the retarded one, but contains the “1” and “2” vertices Keldysh:1964ud . Building on this, the effect of the anisotropy in momentum space generated for instance by the presence of a shear pressure tensor $\pi^{\alpha\beta}$ in photon production in general – including LPM – has been elaborated in Ref. Hauksson:2017udm , paving the way for a future holistic treatment of out-of-equilibrium effects, even if the presence of early medium instabilities requires some care Hauksson:2020wsm .
@@
With all of those aspects in mind and several of them still in flux, it is revealing to compare the spectra of real photons obtained with the non-equilibrium approaches detailed in Refs. Linnyk:2015tha ; Greif:2016jeb ; Oliva:2017pri ; Gale:2021emg and shown in Fig. 3. In each of those calculations, the initial state was treated differently. The AFTm (Abelian Flux Tube model) sets up an initial chromo-electric field which decays and populates the fermion distributions Oliva:2017pri . In PHSD (Parton Hadron String Dynamics) Linnyk:2015tha it is string breaking which initializes the partons, which are evolved with the relativistic transport theory. The BAMPS (Boltzmann Approach to Multiparton Scaterrings) simulations Greif:2016jeb are based on a gluon-dominated initial state, as does the multistage calculation involving KøMPøST discussed in Gale:2021emg , but the evolution times of the classical Yang-Mills equations are a parameter of the model. In BAMPS, the gluon Wigner distributions are sampled from the solution of the classical Yang-Mills and then propagated in the partonic cascade. In the multistage model the energy-momentum tensor is passed from IP-Glasma to KøMPøST and then fed to the hydrodynamic evolution. Clearly all the approaches discussed so far – those with results highlighted in Fig. 3 – differ in the details of their respective implementation even though some of the basic principles are common.

### v2

@@ 1 Introduction
Measurements of the electromagnetic radiation emanating from the collision of strongly interacting particles date back to the very early scattering experiments which predate the establishment of the theory of the nuclear strong interaction, QCD. Owing to the relative weakness of the electromagnetic coupling, $\alpha_{\rm EM}/\alpha_{\rm s}\ll 1$, photons (real and virtual) offered the promise of a clean probe of the reaction dynamics as the mean free paths associated with the electromagnetic interaction exceed typical hadronic length scales. In the modern era ushered in by the availability of high energy hadronic beams, the immense potential of electromagnetic observables became clear, and their status as natural complements to measurements of hadrons was confidently established. To make the discussion specific, the photon production cross section in proton-proton collisions, for example, is calculable in perturbative QCD (pQCD), with the help of factorization Field:1989uq . Consequently, early measurements of electromagnetic radiation at CERN (ISR) and at Fermilab where crucial in the confirmation of the parton model, of QCD itself, and of its perturbative treatment Drell:1970wh ; Owens:1986mp .
To repeat, the basic feature which makes real and virtual photons so appealing in studies of QCD is that they suffer little final state interactions and therefore decouple from the strongly interacting system. They are therefore good probes of the conditions local to their production site, assuming their formation mechanism is under theoretical control. This is also the reason why electromagnetic measurements have accompanied the development of the field now known as relativistic heavy-ion collisions which concerns itself with the formation, study, and characterization of the quark-gluon plasma, an exotic state of the fundamental constituents of QCD and which is also the central theme of the present volume.
@@
Figure 1: The many stages of a heavy-ion collision: (a) Initial state; (b) pre-hydro phase; (c) fluid-dynamical evolution; (d) Hadronization; (e) Final-state hadrons eventually free-streaming to the detectors. These images are not from a single end-to-end calculation, but are simply meant to illustrate the different stages. The pictures have been assembled and collected from different sources which include hydrodynamical simulations Schenke:2010nt and UrQMD calculations Bleicher:1999xi .
@@ 2 Electromagnetic radiation from systems in equilibrium
Some of the early exploratory calculations devoted to estimates of the temperature of the quark-gluon plasma have proposed measuring the spectrum of lepton pairs of invariant mass $M$ in the region $m_{\phi}<M\lesssim m_{J/\psi}$ and extracting an effective temperature from its slope, in analogy with black body radiation Shuryak:1978ij ; Kajantie:1981wg ; Domokos:1980ba ; Kajantie:1986dh . This basic philosophy – using electromagnetic radiation to characterize the QGP – is followed to this day, but with some of the additional sophistication described in this paper.
@@
In a strongly interacting medium at temperature $T=1/\beta$, a starting point to the computation of thermal electromagnetic emission rates is the transition rate between two states, an initial state $i$ and a final state $f$, which can be written as $R_{fi}=\frac{\left|S_{fi}\right|^{2}}{\tau V}$ with $\tau V$ being the proper four-volume. To leading order in the electromagnetic interaction, the interaction matrix element written in terms of the electromagnetic current operator $J^{\mu}$ coupled to the photon field $A^{\mu}$ is $S_{fi}=\langle f|\int\!\!d^{4}xJ^{\mu}(x)A_{\mu}(x)|i\rangle$. The vector field may then be written as $A^{\mu}(x)=\epsilon^{\mu}\left(e^{ik\cdot x}+e^{-ik\cdot x}\right)/\sqrt{2\omega V}$, where $\epsilon^{\mu}$ is a polarization vector and $k^{\mu}=\left(\omega,{\bf k}\right)$. The photon rate is
$R_{fi}=-\frac{g^{\mu\nu}}{2\omega V}(2\pi)^{4}\left[\delta\left(p_{i}+k-p_{f}\right)+\delta\left(p_{i}-k-p_{f}\right)\right]\langle f|J_{\mu}(0)|i\rangle\langle i|J_{\nu}(0)|f\rangle$
where one delta function deals with the absorption process, the other with emission.
Using the equation of motion (or Maxwell’s equations) and the Schwinger-Dyson equation leads to the expressions for the differential emission rate $\Gamma$ (electromagnetic quanta per unit four-volume) of photons and dileptons Feinberg:1976ua ; McLerran:1984ay ; Weldon:1990iw ; Gale:1990pn ; Kapusta:2007xjq :
$\omega\frac{d\Gamma_{\gamma}}{d^{4}k}$ $=$ $-\frac{g^{\mu\nu}}{\left(2\pi\right)^{3}}{\rm Im}\,\Pi_{\mu\nu}(\omega,{\bf k})\frac{1}{e^{\beta\omega}-1}\ \ {\rm(photons)}$
$E_{+}E_{-}\frac{d\Gamma_{\ell\bar{\ell}}}{d^{3}p_{+}d^{3}p_{-}}$ $=$ $\frac{2e^{2}}{\left(2\pi\right)^{6}}\frac{1}{k^{4}}L^{\mu\nu}\,{\rm Im}\Pi_{\mu\nu}(\omega,{\bf k})\frac{1}{e^{\beta\omega}-1}\ \ {\rm(dileptons)}$ (2)
@@
The lepton tensor is (for leptons with rest mass $m_{\ell}$) $L^{\mu\nu}=p_{+}^{\mu}p_{-}^{\nu}+p_{+}^{\nu}p_{-}^{\mu}-g^{\mu\nu}\left(p_{+}\cdot p_{-}+m_{\ell}^{2}\right)$. The photon (dilepton) four-momentum $k^{\mu}$ is light-like for real photons and timelike for dileptons. In these equations, $\Pi^{\mu\nu}$ is the finite-temperature retarded in-medium photon self-energy; it contains the information on how the medium affects the particles eventually assembling into real or virtual photons. Importantly, these equations are valid up to first order in the electromagnetic coupling, but are exact in the strong coupling Gale:1990pn . This formalism is therefore directly amenable to non-perturbative field theoretical calculations, where the photon self-energy can be written as a current-current correlator (the electromagnetic current) Gale:1990pn obtainable by analytically continuing the Euclidean correlator evaluated non-perturbatively on the lattice. For a non-Abelian gauge theory like QCD this is a technically challenging procedure Ghiglieri:2016tvj , but recent progress has been reported Ce:2022fot ; Ali:2024xae .
@@
An alternative approach to the evaluation of electromagnetic radiation production rates is provided by the fact that taking the imaginary part of the photon self-energy is tantamount to putting internal virtual particles on-shell, enabling in principle a formalism based on kinetic theory Weldon:1983jn ; McLerran:1984ay ; Majumder:2001iy . This correspondence is most readily seen in the case of dilepton production where the “simplest” (in terms of loop topology) diagram of the (virtual) photon self-energy is a single fermion loop:
where the dashed line indicates the taking of the imaginary part or colloquially “cuttting the diagram”. The cut diagram then reveals the $q\bar{q}\to\gamma^{*}$ process Gale:1990pn , often calculated using kinetic theory.
In this language, the photon production rate from the reaction $1+2\to 3+\gamma$ is
$\Gamma_{\gamma}$ $=$ ${\cal N}\int\frac{d^{3}p_{1}}{2E_{1}\left(2\pi\right)^{3}}\frac{d^{3}p_{2}}{2E_{2}\left(2\pi\right)^{3}}f_{1}(E_{1})f_{2}(E_{2})\left(2\pi\right)^{4}\delta\left(p_{1}^{\mu}+p_{2}^{\mu}-p_{3}^{\mu}-k^{\mu}\right)$ (3)
$\times\left|{\cal M}\right|^{2}\frac{d^{3}p_{3}}{2E_{3}\left(2\pi\right)^{3}}\frac{d^{3}k}{2\omega\left(2\pi\right)^{3}}\left[1\pm f_{3}(E_{3})\right]$
[…]
@@
The historical development of thermal emission rates from strongly interacting matter as that created in heavy-ion collisions is an interesting story in itself. Calculations of photons and dileptons from the hot QGP involve the interactions between gluons and quarks in a finite-temperature medium. Early kinetic theory calculations provided estimates of photon and dilepton emission from interacting partons Shuryak:1978ij ; Kajantie:1981wg ; Sinha:1983jm ; Hwa:1985xg ; Staadt:1985uc . Writing the production rate of real photons in terms of the photon self-energy with Eq. (2) in mind, the processes at leading order in $\alpha_{s}$: $qg\to q\gamma$ and $q\bar{q}\to g\gamma$ appear at two loops in the loop expansion of the photon self-energy. If the virtual quarks exchanged in those reactions are massless, the emission rate suffers from a logarithmic collinear singularity, cured however by the generation of an in-medium effective fermion mass calculated using the Hard Thermal Loops (HTL) resummation Kapusta:1991qp ; Baier:1991em . The formalism of HTL was established earlier and used to famously address the manifest and vexing gauge dependence that plagued calculations of the gluon damping rate in finite temperature QCD [^1] See Ref. [3] in Ref. Braaten:1989mz ..
@@
Figure 2: Two photon-producing processes that come from taking the imaginary part of a two-loop photon self-energy. They are (left) bremsstrahlung from a quark (or anti-quark), and (right) $q{\bar{q}}$ annihilation with scattering. The gluon propagator is HTL-corrected (see main text).
@@
Further explorations in the topology of the photon self-energy lead to evaluations of the parton bremsstrahlung contribution to the production of real photons. Using an effective theory based on the resummation of HTLs, this channel enters at the two-loop level in that effective theory and taking the imaginary parts produces the processes shown in Fig. 2. Even though the naive parametric power counting of the photon self-energy would have those contributions $\sim g^{4}$, where $g$ is the strong coupling constant, it was found that collinear divergences formally promote the photon rate to $\sim g^{2}$ Aurenche:1998nw . The solution to the puzzling question of how to uncover and consistently sum up all the contributions at a given order of the perturbation expansion was provided with the inclusion of photons generated by the Landau-Pomeranchuk-Migdal effect (LPM), which limits the coherence length of the emitted radiation. The first photon production rates complete at leading order (LO) in the strong coupling were presented in Ref. Arnold:2001ms , and results at NLO were reported later in Ref. Ghiglieri:2013gia  [^2] For values of $\alpha_{s}$ relevant for the phenomenology discussed here, NLO + LO results show an enhancement of $\sim 20\%$ over LO, owing to some partial cancellations in the NLO contributions..
@@
At lower temperatures, the photon rates calculated with the hadronic degrees of freedom relevant to stages (d) and (e) of Fig. 1, have been calculated using interactions modeled with effective chiral hadronic Lagrangians in Refs. Meissner:1987ge ; Kapusta:1991qp ; Song:1993ae ; Turbide:2003si ; Heffernan:2014mla ; Holt:2015cda . Calculations of dilepton production involving composite hadrons also rely on Eq. (2), but the virtual nature of the photon field reveals the vector meson spectral distribution, owing to the coupling of the photon field with vector meson fields: Vector Meson Dominance Gale:1990pn ; Rapp:1999ej .
@@ 2.1 Electromagnetic emission out of equilibrium
To fix ideas, we first consider the emission of real photons; of course, virtual photons are also emitted out of equilibrium. Before discussing the different models addressing photon emission out of equilibrium and just prior to the hydrodynamics phase, it is instructive to discuss the most extreme case: electromagnetic radiation from the very first nucleon-nucleon collisions. Those photons will appear as an irreducible background to any of the other photon sources under consideration, and will typically occupy the highest $p_{T}$ part of the spectrum. As alluded to previously, those are calculable with pQCD, using a procedure that is however not free from uncertainties which will grow as $p_{T}$ decreases: they are associated with scale uncertainties in the nuclear parton density functions (nPDF) Aurenche:2006vj ; Arleo:2011gc ; Klasen:2013mga ; Paakkinen:2017jpo and with the relatively poorly known photon fragmentation function (especially from gluons) Klasen:2014xfa . This is why the pQCD contributions in the low momentum part $p_{T}~\approx~1$ GeV/c of the thermal photon window are mostly extrapolations Paquet:2015lta which emphasize the importance of direct measurements of low transverse momentum photons in pp and pA collisions.
@@
Just prior to, but also in the fluid phase, non-equilibrium aspects will manifest themselves in different ways. In the recent couple of decades, the relativistic heavy-ion program pursuing the characterization of the QGP has revealed the remarkable success of viscous fluid dynamical modeling Gale:2013da . By comparing with hadronic data, it became apparent that the results of the early calculations based on inviscid hydrodynamics could be brought closer to measurements with the inclusion of transport coefficients such as shear and bulk viscosity Gale:2013da ; the electromagnetic emissivity evaluations also needed to be adapted consequently.
@@
Deviations from kinetic equilibrium can be obtained by considering the linearized Boltzmann equation and evaluating the deviations ($\delta f$) of the distribution functions ($f=f_{0}+\delta f$) from their equilibrium form ($f_{0}$), using techniques Denicol pioneered by Chapman and Enskog Chapman , and by Grad Grad . Using that approach, the photoproduction channels in the QGP and in the confined hadronic sector involving $2\to 2$ kinematics can be corrected for the presence of shear and bulk viscosity Dion:2011pp demanded for consistency with the fluid dynamical evolution HeinzSchenke . The electromagnetic emission rates derived thusly can then be implemented in parallel with the fluid dynamical evolution Paquet:2015lta . Part of this general framework for correcting the photon emission rates is discussed in Ref. Dion:2011pp , and a discussion of the $2\to 2$ photon-producing contributions in the QGP (Compton and quark-anti-quark annihilation) appears in Ref. Shen:2014nfa . However, note that not all the photoproduction channels currently used in phenomenology include such out-of-equilibrium corrections: an example is photon production related to the LPM effect.
@@
It is known that formally, a field-theoretic formulation appropriate for non-equilibrium contributions is based on the real-time formalism as expressed in the “1/2” Schwinger1961 ; Keldysh:1964ud and “r/a” bases [^3] The numbers (1/2) label two branches in the complex time plane and (r/a) stands for “retarded” and “advanced”. Keldysh:1964ud . In that language, the thermal photon spectrum is Serreau:2003wr
@@
where the photon polarization tensor is no longer the retarded one, but contains the “1” and “2” vertices Keldysh:1964ud . Building on this, the effect of the anisotropy in momentum space generated for instance by the presence of a shear pressure tensor $\pi^{\alpha\beta}$ in photon production in general – including LPM – has been elaborated in Ref. Hauksson:2017udm , paving the way for an eventual holistic treatment of out-of-equilibrium effects, even if the presence of early medium instabilities requires some care Hauksson:2020wsm .
@@
With all of those aspects in mind and several of them still in flux, it is revealing to compare the spectra of real photons obtained with the non-equilibrium approaches detailed in Refs. Linnyk:2015tha ; Greif:2016jeb ; Oliva:2017pri ; Gale:2021emg and shown in Fig. 3. In each of those calculations, the initial state was treated differently. The AFTm (Abelian Flux Tube model) sets up an initial chromo-electric field which decays and populates the fermion distributions Oliva:2017pri . In PHSD (Parton Hadron String Dynamics) Linnyk:2015tha it is string breaking which initializes the partons, which are evolved with the relativistic transport theory. The BAMPS (Boltzmann Approach to Multiparton Scaterrings) simulations Greif:2016jeb are based on a gluon-dominated initial state, as does the multistage calculation involving KøMPøST discussed in Gale:2021emg , but the evolution times of the classical Yang-Mills equations are a parameter of the model. In BAMPS, the gluon Wigner distributions are sampled from the solution of the classical Yang-Mills and then propagated in the partonic cascade. In the multistage model the energy-momentum tensor is passed from IP-Glasma to KøMPøST and then fed to the hydrodynamic evolution. Clearly all the approaches discussed so far – those with results highlighted in Fig. 3 – differ in the details of their respective implementation even though some of the basic principles are common. The size of the non-equilibrium photon production as compared with the more conventional thermal source can be assessed by examining, for instance, Figure 4.